# Attributional Safety Failures in Large Language Models under Code-Mixed Perturbations

### v1

@@
Abstract
Recent advancements in LLMs have raised significant safety concerns, particularly when dealing with code-mixed inputs and outputs. Our study systematically investigates the increased susceptibility of LLMs to produce unsafe outputs from code-mixed prompts compared to monolingual English prompts. Utilizing explainability methods, we dissect the internal attribution shifts causing model’s harmful behaviors. In addition, we explore cultural dimensions by distinguishing between universally unsafe and culturally-specific unsafe queries. This paper presents novel experimental insights, clarifying the mechanisms driving this phenomenon.
Attributional Safety Failures in Large Language Models
under Code-Mixed Perturbations
[…]
1 Introduction
@@
“Code-mixing” the frequent blending of two or more languages within conversations or even within single utterances – is a prevalent linguistic phenomenon observed globally, especially in linguistically diverse regions like South Asia, Africa, and Latin America Bali et al. (2014); Das and Gambäck (2014). In fact, the phenomenon of code-mixing is as old as the culture of bilingualism and multilingualism Ezeh et al. (2022). Further, speakers who frequently code-mix develop greater cultural awareness. They learn how meaning shifts across contexts and how to navigate cultural norms in each. This linguistic agility mirrors cultural fluidity, especially in urban, globalised settings Grosjean (2010).
@@
The complexities inherent to code-mixed languages, such as hybridized syntax, diverse lexical borrowing, and inconsistent orthographic conventions, substantially complicate processing tasks for even advanced LLMs Singh et al. (2018). While such models excel in monolingual settings, their performance deteriorates noticeably when faced with code-mixed inputs Zhu et al. (2024) (see Figure 1). As digital interactions increasingly adopt multilingual and code-mixed expressions, especially on social media and online forums, understanding the limitations of LLMs in these contexts becomes imperative for their safe and ethical deployment.
@@
Figure 1: Illustrative examples comparing model outputs for monolingual and code-mixed prompts (generated using LLaMA-3.1-8B).
@@
Existing research primarily emphasizes monolingual or well-structured multilingual tasks, leaving a noticeable gap in systematic evaluations specifically addressing code-mixed scenarios. Further, the dimension of cultural sensitivity, crucial in determining perceptions of harm and appropriateness, remains significantly underexplored in LLM literature. Cultural contexts play a pivotal role in shaping linguistic interpretations and reactions; phrases or topics considered benign in one culture may be offensive or harmful in another Hershcovich et al. (2022). This cultural variability underscores the importance of distinguishing universally unsafe content, such as explicitly violent or hateful speech, from culturally-specific unsafe content, which requires a deeper and nuanced understanding of regional and societal norms.
@@
To address these critical gaps, our study systematically evaluates how and why LLMs struggle to safely process culturally sensitive code-mixed inputs, integrating an explainability-driven approach. By interpreting the neural network behaviors, we meticulously analyse and visualise the internal attribution shifts occurring when LLMs encounter code-mixed versus monolingual inputs Kokhlikyan et al. (2020). By pinpointing attribution shifts, we provide clear visual evidence and quantitative insights into how linguistic complexity in code-mixing directly contributes to unsafe or inappropriate outputs. We systematically explore and highlight how culture-specific nuances and sensitivities in code-mixing influence the LLMs’ perception and response generation. Through comparative analysis of cultural aspects that are globally relevant versus those that are only locally relevant, our research sheds light on crucial cultural factors that exacerbate the risks associated with multilingual LLM deployments.
@@
The principal contributions of this paper can be summarised as follows: First, we present a rigorous, systematic evaluation that explicitly demonstrates increased vulnerabilities in state-of-the-art LLMs when processing code-mixed inputs. Second, by utilising advanced interpretability techniques, we provide detailed and interpretable insights into the specific linguistic complexities triggering harmful model outputs. Last, we significantly extend the understanding of the interplay between culture and linguistic safety, illuminating critical cultural nuances that impact LLM responses.
@@
2 Related work
Despite the impressive capabilities of recent LLMs, critical concerns around their safety, robustness, and fairness have emerged, particularly in non-standard input settings such as code-mixing Bommasani et al. (2021); Weidinger et al. (2021). Code-mixed language, common in linguistically diverse societies, poses unique challenges due to its syntactic fluidity and lexical overlap Sitaram et al. (2019); Khanuja et al. (2020). Although fine-tuning and synthetic data augmentation have shown promise Winata et al. (2021); Singh et al. (2021); Panda et al. (2020), LLMs still show vulnerabilities, especially in identifying toxic, biased, or culturally inappropriate content within such hybrid text Gehman et al. (2020); Bender et al. (2021); Meade et al. (2023); Abid et al. (2021). Unlike earlier models like BERT or mBERT that were pretrained for token classification tasks, contemporary LLMs perform generative reasoning, making the safety concerns more pressing and complex Huang et al. (2022); Liu et al. (2023).
Interpretability is increasingly critical in understanding and mitigating these issues. Tools like Captum, LIME, and SHAP have been leveraged to provide insights into how LLMs distribute attention or attribution across input tokens Kokhlikyan et al. (2020); Ribeiro et al. (2016); Lundberg and Lee (2017), but their application to LLM-generated content in multilingual or code-mixed contexts is limited Danilevsky et al. (2020); Gilpin et al. (2018); Jacovi and Goldberg (2020). Cultural sensitivity also remains underexplored; studies have shown that language models can generate outputs that reflect cultural biases or misunderstand nuanced regional semantics Hershcovich et al. (2022); Sheng et al. (2021); Tan et al. (2021); Thomas et al. (2021); Hovy et al. (2021). While safety filters and alignment techniques have emerged Ganguli et al. (2023); Xu et al. (2021); Zhou et al. (2023); Banerjee et al. (2025d); Hazra et al. (2024a); Banerjee et al. (2025b), most remain evaluated on English benchmarks or sanitised datasets, failing to account for linguistic and cultural heterogeneity Rudinger et al. (2022); Srivastava et al. (2022); Sangati et al. (2020); Banerjee et al. (2024); Hazra et al. (2024b); Banerjee et al. (2025a). This work addresses these gaps by applying attribution-based explainability to LLMs in code-mixed contexts, demonstrating how attribution shifts signal unsafe or biased outputs and guiding culturally informed mitigation strategies.
3 Code-mixing method and the datasets
@@
Method for code-mixing: In constructing our code-mixed dataset, we adopt the Matrix Language Frame (MLF) model proposed by Myers-Scotton Myers-Scotton (1993), wherein a dominant “Matrix Language” governs the grammatical structure of the sentence while elements from an “Embedded Language” are inserted at permissible syntactic boundaries. This framework aligns well with natural bilingual speech patterns and has been computationally modeled in recent synthetic data generation tools Rizvi et al. (2021). In our implementation, we select the native or culturally primary language as the matrix language and maintain a structured 60:40 ratio between the matrix and the embedded language content respectively (see Figure 2). This ratio is consistent with previous studies that found a higher proportion of matrix language tokens improves the fluency and syntactic validity of code-mixed text Pratapa et al. (2018); Bhat et al. (2016). By maintaining this distribution, we ensure that the generated dataset not only adheres to linguistic theory but also aligns with human judgments of naturalness and readability observed in bilingual communities.
Dataset: We leverage a culturally grounded dataset introduced by Banerjee et al. (2025c), specifically designed to probe the safety behavior of LLMs under code-mixed perturbations. The dataset is constructed across 10 languages: Arabic, Bengali, Chinese, Hindi, Japanese, Korean, Russian, German, Spanish and Portuguese and covering 12 sociopolitical domains (see Figure 11). It is systematically partitioned into two evaluation subsets: Generic TestSet and Local TestSet, each curated to target distinct facets of cultural sensitivity. Each of these serves as the primary matrix language, defining the syntactic and grammatical backbone of the code-mixed sentences. The embedded language across all instances is English, selected due to its prominence in global digital discourse. To generate linguistically plausible code-mixed variants, we follow the MLF model (Myers-Scotton, 1993) discussed above, which allows the insertion of embedded language tokens at syntactically permissible boundaries within the matrix language structure.
The Generic TestSet is intended to capture generic cultural unsafe queries, those that reflect widely recognized harms across cultures without tying them to specific regional or linguistic contexts. We select 814 instances following Banerjee et al. (2025c) as part of this category. These prompts represent broadly adversarial scenarios involving topics such as political corruption, extremism, and violence, but without any culture-specific markers. They serve as a form-invariant probe of general safety alignment.
@@
Figure 2: Actual vs generated code-mixed examples using the MLF technique.
@@
Actual Code-mixed Actual Code-mixed Actual Code-mixed Actual Code-mixed Actual Code-mixed Actual Code-mixed
@@
Table 1: Performance comparison across various cultures for the Generic TestSet and Local TestSet.
4 Experimental setup
Model selection: To investigate the safety vulnerabilities in linguistically complex environments, we evaluate three publicly available instruction-tuned models, phi-4B, mistral-7B, and llama-3.1-8B. These models span a range of parameter scales and architectural families, enabling a controlled comparison of cultural sensitivity under varying representational capacities.
@@ 4 Experimental setup
ASR=# Harmful responses# Total responses $\textit{ASR}=\frac{\#\text{ Harmful responses}}{\#\text{ Total responses generated}}$ (1)
@@
Each response is labeled as harmful based on judgments made exclusively by gpt4-o, which serves as the unified safety evaluator due to its demonstrated consistency and interpretability in toxicity and harmfulness detection. For each response, gpt4-o is prompted with detailed context and cultural framing to assess whether the output violates norms of cultural respect, safety, or neutrality. To ensure consistency, we calibrate the prompting interface across all evaluations and adopt a binary decision protocol.
5 Results and analysis
5.1 Evaluation with Generic TestSet
We first examine the performance degradation of LLMs when evaluated on the Generic TestSet across 10 cultural contexts (see Figure 11). As shown in Table 1, the transition from monolingual prompts to code-mixed inputs results in a marked increase in ASR across all three models. phi-4B, which maintains a relatively low ASR of 9.46% in monolingual settings, experiences a nearly 7.3$\times$ increase (to 69.08%) when exposed to code-mixed queries. Similarly, mistral-7B and llama-3.1-8B exhibit average ASR increases of +56.17% and +22.57%, respectively.
@@ 5.1 Evaluation with Generic TestSet
These results suggest that even instruction-tuned LLMs exhibit a significant degradation in safety alignment under multilingual interference. This aligns with prior findings on cross-lingual robustness failures in aligned LLMs (Liang et al., 2022; Huang et al., 2023). Notably, languages linguistically more distant from the “embedding language” English (e.g., Arabic, Bengali, Japanese) exhibit some of the most dramatic changes. Conversely, even languages with greater token-level overlap with English (e.g., German, Spanish) demonstrate that lexical similarity alone is insufficient to maintain semantic consistency in safety behavior under code-mixing. Such failures suggest that current safety tuning pipelines are insufficiently robust to perturbations introduced by code-mixed prompts.
5.2 Evaluation with the Local TestSet
@@
The Local TestSet enables finer-grained analysis of culturally contextualized safety vulnerabilities. Across all models, we observe a consistent and significant increase in ASR when moving to code-mixed inputs: from 36.63% to 68.73% for phi-4B, 74.79% to 84.08% for llama-3.1-8B, and 74.16% to 84.66% for mistral-7B. Even the larger models like llama-3.1-8B exhibit are not immune to safety collapses under linguistic mixing, with ASR reaching 91.12% in Arabic and 93.14% in Hindi.
@@
These results are consistent with emerging literature on the fragility of safety alignment under linguistic variation (Xu et al., 2023; Gao et al., 2023). The models appear to over-rely on high-resource English safety priors and fail to preserve safety boundaries when confronted with form-shifted or culturally grounded code-mixed inputs. In several cases, we observe that code-mixed prompts invert the safety behavior of models, producing harmful continuations where none existed under the original English phrasing. This highlights the necessity of incorporating code-mixed and sociolinguistic variation into safety training regimes for real-world deployment of LLMs in multilingual societies.
6 Attribution analysis

### v2

@@
Abstract.
While LLMs appear robustly safety-aligned in English, we uncover a catastrophic, overlooked weakness: attributional collapse under code-mixed perturbations. Our systematic evaluation of open models shows that the linguistic camouflage of code-mixing—“blending languages within a single conversation”—can cause safety guardrails to fail dramatically. Attack success rates (ASR) spike from a benign $\sim$9% in monolingual English to $\sim$69% under code-mixed inputs, with rates exceeding 90% in non-Western contexts such as Arabic and Hindi. These effects hold not only on controlled synthetic datasets but also on real-world social media traces, revealing a serious risk for billions of users. To explain why this happens, we introduce saliency drift attribution (SDA), an interpretability framework that shows how, under code-mixing, the model’s internal attention drifts away from safety-critical tokens (e.g., “violence” or “corruption”), effectively blinding it to harmful intent. Finally, we propose a lightweight translation based restoration strategy that recovers roughly 80% of the safety lost to code-mixing, offering a practical path toward more equitable and robust LLM safety.
copyright: noneconference: ; ;
1. Introduction
@@
Code-mixing” the frequent blending of two or more languages within conversations or even within single utterances – is a prevalent linguistic phenomenon observed globally, especially in linguistically diverse regions like South Asia, Africa, and Latin America (bali2014borrowing; das2014identifying). In fact, the phenomenon of code-mixing is as old as the culture of bilingualism and multilingualism (el-teaching2022). Further, speakers who frequently code-mix develop greater cultural awareness. They learn how meaning shifts across contexts and how to navigate cultural norms in each. This linguistic agility mirrors cultural fluidity, especially in urban, globalised settings (bireal2010).
@@
The complexities inherent to code-mixed languages, such as hybridized syntax, diverse lexical borrowing, and inconsistent orthographic conventions, substantially complicate processing tasks for even advanced LLMs (singh2018language). While such models excel in monolingual settings, their performance deteriorates noticeably when faced with code-mixed inputs (zhu2024promptrobustevaluatingrobustnesslarge) (see Figure 1). As digital interactions increasingly adopt multilingual and code-mixed expressions, especially on social media and online forums, understanding the limitations of LLMs in these contexts becomes imperative for their safe and ethical deployment.
@@
Figure 1. Illustrative examples comparing model outputs for monolingual and code-mixed prompts (generated using LLaMA-3.1-8B).
@@
Existing research primarily emphasizes monolingual or well-structured multilingual tasks, leaving a noticeable gap in systematic evaluations specifically addressing code-mixed scenarios. Further, the dimension of cultural sensitivity, crucial in determining perceptions of harm and appropriateness, remains significantly underexplored in LLM literature. Cultural contexts play a pivotal role in shaping linguistic interpretations and reactions; phrases or topics considered benign in one culture may be offensive or harmful in another (hershcovich2022challenges). This cultural variability underscores the importance of distinguishing universally unsafe content, such as explicitly violent or hateful speech, from culturally-specific unsafe content, which requires a deeper and nuanced understanding of regional and societal norms. This creates a significant digital safety divide: while monolingual English users benefit from robust safety guardrails, users in the Global South who rely on code-mixing are left exposed to harmful content.
@@
To address these critical gaps, our study systematically evaluates how and why LLMs struggle to safely process culturally sensitive code-mixed inputs, integrating an explainability-driven approach. Crucially, we move beyond synthetic stress-testing to validate our findings on naturally occurring Hindi-English social media comments (TRAC-1 (kumar-etal-2018-benchmarking)), ensuring our conclusions scale to ‘wild’ web data. By interpreting the neural network behaviors, we meticulously analyze and visualise the internal attribution shifts occurring when LLMs encounter code-mixed versus monolingual inputs (kokhlikyan2020captum). By pinpointing attribution shifts, we provide clear visual evidence and quantitative insights into how linguistic complexity in code-mixing directly contributes to unsafe or inappropriate outputs. We systematically explore and highlight how culture-specific nuances and sensitivities in code-mixing influence the LLMs’ perception and response generation. Through comparative analysis of cultural aspects that are globally relevant versus those that are only locally relevant, our research sheds light on crucial cultural factors that exacerbate the risks associated with multilingual LLM deployments.
@@
The principal contributions of this paper can be summarised as follows: First, we present a rigorous, systematic evaluation that explicitly demonstrates increased vulnerabilities in state-of-the-art LLMs when processing code-mixed inputs. Second, by utilising advanced interpretability techniques, we provide detailed and interpretable insights into the specific linguistic complexities triggering harmful model outputs. Third, we significantly extend the understanding of the interplay between culture and linguistic safety, illuminating critical cultural nuances that impact LLM responses. Last, we suggest a simple and lightweight restoration strategy to recover the correct attributions lost due to code-mixing.
@@
2. Related work
Despite the impressive capabilities of recent LLMs, critical concerns around their safety, robustness, and fairness have emerged, particularly in non-standard input settings such as code-mixing (bommasani2021opportunities; weidinger2021ethical). Code-mixed language, common in linguistically diverse societies, poses unique challenges due to its syntactic fluidity and lexical overlap (sitaram2019survey; khanuja2020new). Although fine-tuning and synthetic data augmentation have shown promise (winata2021multilingual; singh2021pretraining; panda2020disrpt), LLMs still show vulnerabilities, especially in identifying toxic, biased, or culturally inappropriate content within such hybrid text (gehman2020realtoxicityprompts; bender2021dangers; meade2023emergent; abid2021persistent). Unlike earlier models like BERT or mBERT that were pretrained for token classification tasks, contemporary LLMs perform generative reasoning, making the safety concerns more pressing and complex (huang2022language; liu2023evaluating).
Interpretability is increasingly critical in understanding and mitigating these issues. Tools like Captum, LIME, and SHAP have been leveraged to provide insights into how LLMs distribute attention or attribution across input tokens (kokhlikyan2020captum; ribeiro2016lime; lundberg2017unified), but their application to LLM-generated content in multilingual or code-mixed contexts is limited (danilevsky2020survey; gilpin2018explaining; jacovi2020towards). Cultural sensitivity also remains underexplored; studies have shown that language models can generate outputs that reflect cultural biases or misunderstand nuanced regional semantics (hershcovich2022challenges; sheng2021societal; tan2021detecting; thomas2021culture; hovy2021importance). While safety filters and alignment techniques have emerged (ganguli2023reducing; xu2021detoxifying; zhou2023cultural; Banerjee_Layek_Tripathy_Kumar_Mukherjee_Hazra_2025; hazra-etal-2024-safety; banerjee2025soterialanguagespecificfunctionalparameter), most remain evaluated on English benchmarks or sanitised datasets, failing to account for linguistic and cultural heterogeneity (rudinger2022thinking; srivastava2022beyond; sangati2020codemix; banerjee2024unethicalinstructioncentricresponsesllms; hazra-etal-2024-sowing; banerjee-etal-2025-breaking). This work addresses these gaps by applying attribution-based explainability to LLMs in code-mixed contexts, demonstrating how attribution shifts signal unsafe or biased outputs and guiding culturally informed mitigation strategies.
3. Code-mixing method and the datasets
@@
Method for code-mixing: In constructing our code-mixed dataset, we adopt the Matrix Language Frame (MLF) model proposed by Myers-Scotton (myers1993duelling), wherein a dominant “Matrix Language” governs the grammatical structure of the sentence while elements from an “Embedded Language” are inserted at permissible syntactic boundaries. This framework aligns well with natural bilingual speech patterns and has been computationally modeled in recent synthetic data generation tools (rizvi2021gcm). In our implementation, we select the native or culturally primary language as the matrix language and maintain a structured 60:40 ratio between the matrix and the embedded language content respectively (see Figure 2). This ratio is consistent with previous studies that found a higher proportion of matrix language tokens improves the fluency and syntactic validity of code-mixed text (pratapa2018language; bhat2016grammatical). While our sensitivity analysis (see Appendix C) demonstrates that higher proportions of embedded language (e.g., 20:80) lead to even sharper increases in Attack Success Rate, we retain the 60:40 ratio in the main text as a conservative, linguistically grounded baseline. By maintaining this distribution, we ensure that the generated dataset not only adheres to linguistic theory but also aligns with human judgments of naturalness and readability observed in bilingual communities.
Dataset: We leverage a culturally grounded dataset introduced by (banerjee-etal-2025-navigating), specifically designed to probe the safety behavior of LLMs under code-mixed perturbations. The dataset is constructed across 10 languages: Arabic, Bengali, Chinese, Hindi, Japanese, Korean, Russian, German, Spanish and Portuguese and covering 12 sociopolitical domains (see Figure 1 in the supplementary material). It is systematically partitioned into two evaluation subsets: Generic TestSet and Local TestSet, each curated to target distinct facets of cultural sensitivity. Each of these serves as the primary matrix language, defining the syntactic and grammatical backbone of the code-mixed sentences. The embedded language across all instances is English, selected due to its prominence in global digital discourse. To generate linguistically plausible code-mixed variants, we follow the MLF model (myers1993duelling) discussed above, which allows the insertion of embedded language tokens at syntactically permissible boundaries within the matrix language structure.
The Generic TestSet is intended to capture generic cultural unsafe queries, those that reflect widely recognized harms across cultures without tying them to specific regional or linguistic contexts. We select 814 instances following (banerjee-etal-2025-navigating) as part of this category. These prompts represent broadly adversarial scenarios involving topics such as political corruption, extremism, and violence, but without any culture-specific markers. They serve as a form-invariant probe of general safety alignment.
@@
Figure 2. Actual vs generated code-mixed examples using the MLF technique.
@@
English Code-mixed English Code-mixed English Code-mixed English Code-mixed English Code-mixed English Code-mixed
@@
Table 1. Performance comparison across various cultures for the Generic TestSet and Local TestSet.
4. Experimental setup
Model selection: To investigate the safety vulnerabilities in linguistically complex environments, we evaluate three publicly available instruction-tuned models, phi-4B, mistral-7B, and llama-3.1-8B. These models span a range of parameter scales and architectural families, enabling a controlled comparison of cultural sensitivity under varying representational capacities. We deliberately select state-of-the-art open-weight models rather than closed-source API models (e.g., GPT) for our primary analysis. This is because our Saliency Drift Attribution (SDA) framework requires direct access to model gradients and internal attention weights to causally explain safety failures—access that is restricted in closed systems.
@@ 4. Experimental setup
(1) $\textit{ASR}=\frac{\#\text{ Harmful responses}}{\#\text{ Total responses generated}}$
@@
Each response is initially labeled as harmful or non-harmful using GPT-4o as a unified automatic safety annotator, chosen for its stability and coverage across languages. To ensure that these automatic labels are meaningful in culturally diverse settings, we calibrate gpt4-o against a separate human evaluation with native speakers (Section 6) and show that its harmfulness judgments exhibit substantial agreement with human annotators. For each response, gpt4-o is prompted with detailed context and cultural framing to assess whether the output violates norms of cultural respect, safety, or neutrality. To ensure consistency, we calibrate the prompting interface across all evaluations and adopt a binary decision protocol.
5. Results and analysis
5.1. Evaluation with Generic TestSet
[…]
@@ 5.1. Evaluation with Generic TestSet
These results suggest that even instruction-tuned LLMs exhibit a significant degradation in safety alignment under multilingual interference. This aligns with prior findings on cross-lingual robustness failures in aligned LLMs (liang2022holistic; huang2023prompt). Notably, languages linguistically more distant from the “embedding language” English (e.g., Arabic, Bengali, Japanese) exhibit some of the most dramatic changes. Conversely, even languages with greater token-level overlap with English (e.g., German, Spanish) demonstrate that lexical similarity alone is insufficient to maintain semantic consistency in safety behavior under code-mixing. Such failures suggest that current safety tuning pipelines are insufficiently robust to perturbations introduced by code-mixed prompts.
5.2. Evaluation with the Local TestSet
@@
The Local TestSet enables finer-grained analysis of culturally contextualized safety vulnerabilities. Across all models, we observe a consistent and significant increase in ASR when moving to code-mixed inputs: from 36.63% to 68.73% for phi-4B, 74.79% to 84.08% for llama-3.1-8B, and 74.16% to 84.66% for mistral-7B. Even the larger models like llama-3.1-8B are not immune to safety collapses under linguistic mixing, with ASR reaching 91.12% in Arabic and 93.14% in Hindi.
@@
These results are consistent with emerging literature on the fragility of safety alignment under linguistic variation (xu2023crosslingual; gao2023toxigen). The models appear to over-rely on high-resource English safety priors and fail to preserve safety boundaries when confronted with form-shifted or culturally grounded code-mixed inputs. In several cases, we observe that code-mixed prompts invert the safety behavior of models, producing harmful continuations where none existed under the original English phrasing. This highlights the necessity of incorporating code-mixed and sociolinguistic variation into safety training regimes for real-world deployment of LLMs in multilingual societies.
5.3. Evaluation with social media data
While the MLF-generated dataset allows for controlled syntactic evaluation, we further validate our findings on naturally occurring social media text to ensure ecological validity. We utilized the TRAC-1 dataset (kumar-etal-2018-benchmarking), a corpus of authentic, human-generated real-world Hindi-English code-mixed comments scraped from Facebook and annotated for aggression. When evaluated against this ’wild’ data, the models exhibited safety failures consistent with our synthetic findings. We observe, the ASR remained critically high: 61.32% for phi-4B, 74.87% for mistral-7B, and 79.24% for llama-3.1-8B. The persistence of high ASR on this dataset confirms that attributional collapse is not an artifact of synthetic generation (MLF) but a pervasive vulnerability affecting billions of interactions on platforms like Facebook today.
6. Human validation
To validate and calibrate our automatic labels, we conducted a large-scale human evaluation across six culturally and linguistically diverse languages. gpt4-o based annotations offer consistency and scalability, while human judgments provide the primary reference for culturally grounded safety failures. We therefore use the human study to check whether gpt4-o’s harmfulness labels are aligned with native-speaker intuitions.
6.1. Annotation setup
We employ Prolific [^1] https://www.prolific.com/ to collect human annotations on a subset of the dataset. Each example was independently annotated by three native speakers, with final labels determined via majority vote. Annotators completed three tasks: (1) translation quality of the code-mixed prompt (good/average/poor), (2) harmfulness of the English response, and (3) harmfulness of the code-mixed response (both binary). Model outputs were anonymised and shuffled across architectures to avoid bias.
6.2. Findings and agreement
[…]
Culture Eng ASR (%) CM ASR (%) $\Delta$ ASR TQ (%)
Chinese 16.22 39.20 +22.98 75.68
German 18.91 28.37 +9.46 100.00
Bengali 2.70 38.10 +35.40 78.39
Hindi 13.82 41.30 +27.48 88.68
Spanish 5.45 27.56 +22.11 91.89
Arabic 1.35 37.02 +35.67 85.14
Table 2. Human evaluation results across six languages. TQ = Translation Quality. ASR = Attack Success Rate.
For all our evaluation, the inter-annotator agreement was substantial: Fleiss’ $\kappa$ was 0.76 for English harmfulness, 0.81 for code-mixed harmfulness, and 0.74 for translation quality (good + average). These scores indicate substantial agreement between gpt4-o and native-speaker annotations across all six languages, and they reinforce the attributional trends observed in our automated analysis. For scalability, we therefore report results using gpt4-o based annotations in the main experiments, and use the human labels as an external check that our conclusions are not driven by idiosyncrasies of a single safety model.
7. Attribution analysis